# Physics of droplet regulation in biological cells

### v2

@@ 2.1.4  Interfaces and phase morphologies
In the absence of chemical reactions ($s=0$), the dynamics described by equation (2.23) result in equilibrium states that minimise the free energy $F$ given by equation (2.22). Figure 2.2A shows that the equilibrium profile $\phi_{\mathrm{eq}}(x)$ smoothly interpolates between the fractions $\phi^{(0)}_{\mathrm{out}}$ and $\phi^{(0)}_{\mathrm{in}}$ of the two co-existing phases. Using a coordinate $x$ perpendicular to the interface placed at $x=0$, the profile can be approximated as
@@
with a typical interfacial width $0pt$ of
$0pt$ $\approx\frac{1}{\phi^{(0)}_{\mathrm{in}}-\phi^{(0)}_{\mathrm{out}}}\,\sqrt{\frac{6\nu\kappa}{k_{\mathrm{B}}T}}$ (2.25)
for the symmetric case $\Omega=1$; see appendix C.
@@
see appendix C. The surface energy $\gamma$ increases with larger gradient parameter $\kappa$, smaller molecular volume $\nu$, and larger segregation (larger $\phi^{(0)}_{\mathrm{out}}-\phi^{(0)}_{\mathrm{in}}$). It can often be measured experimentally [129] or determined from molecular simulations [130] and thus provide a useful parametrisation of interfacial physics.
The thin-interface approximation simplifies the analysis of equilibrium states in binary liquids: material conservation fixes the total volume of the dilute and dense phases, so their shape is the only degree of freedom. The total free energy thus comprises a constant term covering the bulk contributions and an interfacial term, $\gamma A$, where $A$ is the total interfacial area. Equilibrium states thus minimise $A$, which implies zero mean curvature of the interfaces [131]. Consequently, surface energies can be interpreted as a mechanical surface tension [132], explaining why the minority phase favours spherical shapes. Conversely, if both phases occupy similar volumes, bicontinuous structures, such as cylinders and flat lamella, can emerge; see figure 2.2. These structures predominate in a region of phase space that overlaps significantly with the spinodal region; see figure 2.1E,F. However, these two phenomena are distinct, so droplets can for instance emerge from spinodal decomposition. Bicontinuous structures are rarely observed in biological cells, and so we exclusively focus on droplets in this review.
@@ 2.2.1  Local equilibrium at the interface
where $d$ is the dimension of space. These equations reduce to equations (2.5) if surface tension is negligible, $\gamma=0$. However, mechanical equilibrium as expressed in equation (2.30b) reveals that $\gamma$ generates an extra pressure $P_{\gamma}=(d-1)\gamma/R$, which is known as the Laplace pressure. Equations (2.30) are solved by the coexisting equilibrium fractions $\phi^{\mathrm{eq}}_{\mathrm{in}}$ and $\phi^{\mathrm{eq}}_{\mathrm{out}}$, which depend on the droplet radius $R$ and only coincide with $\phi^{(0)}_{\mathrm{in}}$ and $\phi^{(0)}_{\mathrm{out}}$ for flat interfaces ($R\rightarrow\infty$). A pressure increase $\Delta P$ inside droplets generally implies larger chemical potentials, and thus larger $\phi^{\mathrm{eq}}_{\mathrm{in}}$ and $\phi^{\mathrm{eq}}_{\mathrm{out}}$. In the simple case of an incompressible droplet originating from strong phase separation, the fraction inside will hardly change, $\phi^{\mathrm{eq}}_{\mathrm{in}}\approx\phi^{(0)}_{\mathrm{in}}$, whereas the volume fraction outside approximately reads [134]
@@ 2.2.2  Classical nucleation theory
The free energy of an isolated droplet, given by equation (2.28), captures the trade-off between bulk terms that promote phase separation and a surface term that suppresses it. This trade-off becomes particularly important at the initial phase of droplet formation, where a small droplets is embedded in a dilute phase characterised by a fraction $\phi_{\infty}$ away from the droplet. The free energy difference between such a small droplet and the corresponding homogeneous phase without a droplet can be expressed as [136, 76, 137]
@@
where $\Delta\!f=\Pi(\phi^{\mathrm{eq}}_{\mathrm{in}})-\Pi(\phi_{\infty})-[\bar{\mu}(\phi^{\mathrm{eq}}_{\mathrm{in}})-\bar{\mu}(\phi_{\infty})]c_{\mathrm{in}}$ is the driving force for phase separation, and $c_{\mathrm{in}}=\phi^{\mathrm{eq}}_{\mathrm{in}}/(\nu\Omega)$. Note that $\Delta\!f$ is constant in the typical case when phase separation is strong ($\phi^{\mathrm{eq}}_{\mathrm{in}}\approx\phi^{(0)}_{\mathrm{in}}$) and the dilute phase is large (constant $\phi_{\infty}$). In the simple case where the solvent is very mobile (implying pressures equilibrate quickly), and the dilute phase can be described as an ideal solution, $\bar{\mu}_{\infty}\approx k_{\mathrm{B}}T\ln(\phi_{\infty})$, we use equation (2.30a) to find $\Delta\!f\approx k_{\mathrm{B}}Tc_{\mathrm{in}}\ln(\phi_{\infty}/\phi^{\mathrm{eq}}_{\mathrm{out}})$. Supersaturated solutions ($\phi_{\infty}>\phi^{\mathrm{eq}}_{\mathrm{out}}$) thus lower $\Delta F$ and facilitate droplet formation.
Figure 2.3: Free energy $\Delta\!F$ of a small droplet of radius $R$ relative to the homogeneous phase in $2$ and $3$ dimensions; see equation (2.33). The maximum marks the critical size $R_{\mathrm{nucl}}$ and the nucleation barrier $\Delta\!F_{\mathrm{nucl}}$.
@@ 2.2.3  Growth of a single droplet
Once a droplet has nucleated, it can grow by taking up material from its surroundings or by creating material inside itself via chemical reactions. For simplicity, we postpone the discussion of reactions to section 5 and first focus on the case without reactions ($s=0$). Growth of droplets of any shape can be described by the normal velocity $v_{\perp}$ of the droplet interface, which is given by [139, 23]
@@
As a simple example, we consider a spherical droplet embedded in a large system. Without chemical reactions, the droplet is essentially homogeneous, implying $\boldsymbol{j}_{\mathrm{in}}=0$. Conversely, the fraction $\phi_{\infty}$ far away from the droplet might deviate from its equilibrium fraction $\phi^{\mathrm{eq}}_{\mathrm{out}}$, resulting in diffusive fluxes around the droplet. Solving equation (2.37) at stationary state in three dimensions, we find ${\boldsymbol{j}_{\mathrm{out}}=(\phi^{\mathrm{eq}}_{\mathrm{out}}-\phi_{\infty})D_{\mathrm{out}}R^{-1}}$. The droplet growth rate following from equation (2.36) then reads [140]
@@
which gives an explicit equation for $V$ in the limit $\phi^{\mathrm{eq}}_{\mathrm{out}}\approx\phi^{(0)}_{\mathrm{out}}$. This equation illustrates that droplet size is controlled by material conservation (equation 2.29), and that no droplets exist for volume fractions left of the binodal line ($\bar{\phi}<\phi^{(0)}_{\mathrm{out}}$).
@@ 2.3.1  Global parameters
Droplet size is directly influenced by the amount of droplet material in the system, $\bar{\phi}V_{\mathrm{sys}}$; see equation (2.42). Cells can thus influence whether droplets form and how large they get by producing and degrading the relevant droplet material. Droplet size also depends on the coexisting volume fractions $\phi^{(0)}_{\mathrm{out}}$ and $\phi^{(0)}_{\mathrm{in}}$, which in turn depend on the interaction $\chi$ and the molecular size $\Omega$; see equation (2.21). These aspects can be influenced by modifying the involved molecules, either by post-translational modifications [35, 53] or by multimerisation [165], which can induce phase separation in vitro [166] and in vivo [13, 167]. The interactions between molecules are also sensitive to environmental influences [68], including temperature [168, 169], pH [170], osmotic conditions [171], crowding conditions [65], and CO2 concentration [172]. Cells can exploit the phase transition associated with changes in these environmental parameters to deploy appropriate responses e.g., to control flowering in plants [173]. There is also evidence that evolution adapts genetic sequences to different temperature niches to control this behaviour [174].
@@ 2.3.2  Spatial gradients
The sensitivity of phase separation to environmental parameters provides exquisite control over droplet positions when external gradients are involved. For instance, compositional gradients can bias droplets toward one side of a system [110, 175, 176, 177], for both passive [141, 178] and active droplets [179], in a process akin to diffusiophoresis [180]. In sufficiently large cells, gravity can also become important, so that droplets sediment due to density mismatch [78]. Finally, artificially induced temperature gradients can position droplets [168], although such gradients are likely negligible in most cells [98].
@@ 3 Beyond simple liquid droplets
Cells contain an endless variety of biomolecules. For example, there are an estimated $10^{5}$ different proteins [189] present in the cell. It is thus not surprising that condensates typically consist of many different types of molecule [190, 191, 192, 54], many of which have complex individual properties themselves. Molecules can also form partially ordered structures within droplets, such as filaments [193] and porous networks [194]. The complex compositional features resulting from this mixture of biomolecules affects both the bulk and interfacial properties of droplets, and also enables the coexistence of multiple different types of droplet within a cell.
@@ 3.1 Complex molecular interactions
Condensates contain a huge variety of components, including water, ions, small peptides, proteins and flexible polymers such as RNA and DNA [59, 22]. Since it is challenging to directly study interactions between these molecules in their native environment, much experimental work has thus far relied on in vitro reconstitution [195, 196, 126, 37]. Here, it is crucial to exploit label-free techniques [197, 198] since fluorescent labels typically strongly affect droplet condensation [198, 199]. One key insight from these approaches is that often a few scaffold components control the formation of condensates, while client molecules partition into them [200, 201, 202, 203, 204]. However, this dichotomy is likely an oversimplification and it will be interesting to more rigorously quantify the contributions of individual components, for example by using the recently developed dominance metric [205], which can uncover basic interaction motifs between components [206].
It is clear that biomolecular phase separation involves many different kinds of interactions between molecules [207]. To distill the essential features of these interactions, it is often helpful to coarse-grain particular microscopic models [88, 208]. For instance, many biomolecules can be conceptualised as associative polymers [209, 210, 211, 212, 81, 207] and described by sticker-and-spacer models [209, 213, 82, 22, 214], where flexible spacers link stickers with specific interactions. However, there are also many other approaches to modelling the complex interplay of biomolecules that leads to phase separation [22, 85, 59, 215]. Such models allow us to elucidate how particular molecular features, like the amino acid sequence of proteins, affect condensate properties [216, 217, 218, 219], and can explain why dilute phases often exhibit surprisingly large clusters [220, 221].
@@ 3.2 Complex material properties
The complex nature of the constituent biomolecules also results in condensates exhibiting complex rheological properties that cannot be captured by simple (Newtonian) droplet models [129], due to long, flexible polymers deforming, entangling, and interacting with each other. Cells exploit these complex material properties to carry out function: for example, centrosomes are able to grow rapidly in a fluid-like manner, but also exhibit solid-like properties to support large stresses during cell division [228].
@@ 3.2.1  Internal stresses
For a single-component Newtonian fluid, the total stress can be decomposed into contributions arising from hydrostatic pressure, $p$, and viscous dissipation [115]. Multi-component fluids additionally induce an ‘equilibrium stress’, $\boldsymbol{\sigma}_{\textrm{eq}}$, which captures osmotic and interfacial effects. Finally, many biological fluids exhibit a complex rheology that results in an elastic contribution to the stress, denoted by $\boldsymbol{\sigma}_{\textrm{el}}$. The overall stress tensor $\boldsymbol{\sigma}$ then comprises the sum of each of these different contributions,
@@ 3.2.2  Viscoelasticity
Viscoelastic materials can exhibit either liquid-like or solid-like responses to deformation depending on the specific loading conditions. On long timescales (slow deformations), they flow like a liquid, but on shorter timescales (fast deformations), they resist deformation, like an elastic solid. Recent micro-rheological experiments have shown viscoelastic behaviour in a variety of different condensates [129, 233, 234], in which stress relaxation occurs over biologically relevant timescales.
@@
The viscoelastic properties of condensates may change over time, in a process known as ageing [234, 242, 243, 244, 233]. During ageing, biomolecules increasingly attach to binding sites with a higher affinity [67, 245]. This behaviour can be explained by considering a microscopic stickers-and-spacers model which captures competing entropic and elastic effects [246]. As such, aged condensates are less able to readily deform, resulting in a larger relaxation time and a more solid-like rheology. Aged condensates take longer to recover their shape after deformation and show suppressed ability to coalesce [247]. It is also possible for aged condensates to form complex droplet architectures, in which different parts of the droplet display distinct mechanical properties [248]. Note that the storage and loss moduli of aged condensates are often self-similar such that they collapse when scaled by their (time-dependent) characteristic frequency $\omega_{\mathrm{c}}$ [234, 249, 245].
@@ 3.2.3  Gelation
Liquid droplets are characterised by their dynamic, amorphous structure and ability to dissolve in a reversible manner. It is possible, however, for condensates to undergo a phase transition to a gel-like state [13, 250, 227, 251]. This transition results in kinetic arrest, in which condensates maintain their amorphous structure but lose their dynamic properties [252, 253]. Surprisingly, gelation can lead to enhanced ripening relative to liquid droplets [254]. A further transition can result in the formation of irreversible fibrous aggregates, which are linked to the development of various neurodegenerative conditions [27, 12]. Recent work has shown that gelation typically initiates at the interface between the droplet and its surroundings, propagating inwards, such that droplets transiently form a core–shell geometry with a liquid phase surrounded by a gel phase [255]. Similarly, fast changes in composition can result in the formation of kinetically arrested double emulsion structures within droplets[256]. Gelation and aggregation are enhanced in droplets compared to dilute proteins as a result of the increased protein concentration [257, 258]. Gelation can also be driven by shear flows which help to align protein molecules into a kinematically-arrested orientation [259].

### v3

@@ 2.1.4 Interfaces and phase morphologies
In the absence of chemical reactions ($s=0$), the dynamics described by equation (2.23) result in equilibrium states that minimise the free energy $F$ given by equation (2.22). Figure 2.2A shows that the corresponding equilibrium profile $\phi_{\mathrm{eq}}(x)$ smoothly interpolates between the fractions $\phi^{(0)}_{\mathrm{out}}$ and $\phi^{(0)}_{\mathrm{in}}$ of the two co-existing phases. Denoting by $x$ a coordinate perpendicular to the interface positioned at $x=0$, the profile can be approximted as
@@
where $0pt$ denotes the interfacial width. This expression is exact for a simpler polynomial free energy [134, 24]. For our case and symmetric molecular sizes ($\Omega=1$), the interfacial width $0pt$ is approximately (appendix C)
$0pt$ $\approx\frac{1}{\phi^{(0)}_{\mathrm{in}}-\phi^{(0)}_{\mathrm{out}}}\,\sqrt{\frac{6\nu\kappa}{k_{\mathrm{B}}T}}\;,$ (2.25)
and the sigmoidal profile given by equation (2.24) provides a good approximation.
@@
[…]
The thin-interface approximation simplifies the analysis of equilibrium states in binary liquids: material conservation fixes the total volume of the dilute and dense phases, so their shape is the only degree of freedom. The total free energy thus comprises a constant term covering the bulk contributions and an interfacial term, $\gamma A$, where $A$ is the total interfacial area. Equilibrium states thus minimise $A$, which implies constant mean curvature of interfaces [138]. Consequently, surface energies can be interpreted as a mechanical surface tension [139], explaining why the minority phase favours spherical shapes. Conversely, if both phases occupy similar volumes, bicontinuous structures, such as cylinders and flat lamella, can emerge; see figure 2.2. These structures predominate in a region of phase space that overlaps significantly with the spinodal region; see figure 2.1E,F. However, these two phenomena are distinct, so droplets can for instance emerge from spinodal decomposition. Bicontinuous structures are rarely observed in biological cells, and so we exclusively focus on droplets in this review.
@@ 2.2.1 Local equilibrium at the interface
where $d$ is the dimension of space. These equations reduce to equations (2.5) if surface tension is negligible, $\gamma=0$. However, mechanical equilibrium as expressed in equation (2.30b) reveals that $\gamma$ generates an extra pressure $P_{\gamma}=(d-1)\gamma/R$, which is known as Laplace pressure. Equations (2.30) are solved by the coexisting equilibrium fractions $\phi^{\mathrm{eq}}_{\mathrm{in}}$ and $\phi^{\mathrm{eq}}_{\mathrm{out}}$, which depend on the droplet radius $R$ and only coincide with $\phi^{(0)}_{\mathrm{in}}$ and $\phi^{(0)}_{\mathrm{out}}$ for flat interfaces ($R\rightarrow\infty$). A pressure increase $\Delta P$ inside droplets generally implies larger chemical potentials, and thus larger $\phi^{\mathrm{eq}}_{\mathrm{in}}$ and $\phi^{\mathrm{eq}}_{\mathrm{out}}$. In the simple case of an incompressible droplet originating from strong phase separation, the fraction inside will hardly change, $\phi^{\mathrm{eq}}_{\mathrm{in}}\approx\phi^{(0)}_{\mathrm{in}}$, whereas the volume fraction outside approximately reads [141]
@@ 2.2.2 Classical nucleation theory
The free energy of an isolated droplet, given by equation (2.28), captures the trade-off between bulk terms that promote phase separation and a surface term that suppresses it. This trade-off becomes particularly important at the initial phase of droplet formation, where a small droplet is embedded in a dilute phase characterised by a fraction $\phi_{\infty}$ away from the droplet. The free energy difference between such a small droplet and the corresponding homogeneous phase without a droplet can be expressed as [80, 143]
@@
[…]
Figure 2.3: Free energy $\Delta\!F$ of a small droplet of radius $R$ relative to the homogeneous phase in $2$ and 33 dimensions; see equation (2.33). The maximum marks the critical size $R_{\mathrm{nucl}}$ and the nucleation barrier $\Delta\!F_{\mathrm{nucl}}$.
@@ 2.2.3 Growth of a single droplet
Once a droplet has nucleated, it can grow by taking up material from its surroundings or by creating material inside itself via chemical reactions. For simplicity, we postpone the discussion of reactions to section 5 and first focus on the case without reactions ($s=0$). Growth of droplets of any shape can be described by the normal velocity $v_{\perp}$ of the droplet interface, which is given by [146, 147, 148, 24]
@@
As a simple example, we consider a spherical droplet embedded in a large system. Without chemical reactions, the droplet is essentially homogeneous, implying $\boldsymbol{j}_{\mathrm{in}}=0$. Conversely, the fraction $\phi_{\infty}$ far away from the droplet might deviate from its equilibrium fraction $\phi^{\mathrm{eq}}_{\mathrm{out}}$, resulting in diffusive fluxes around the droplet. Solving equation (2.37) at stationary state in three dimensions, we find ${\boldsymbol{j}_{\mathrm{out}}=(\phi^{\mathrm{eq}}_{\mathrm{out}}-\phi_{\infty})D_{\mathrm{out}}R^{-1}}\boldsymbol{e}_{r}$. The droplet growth rate following from equation (2.36) then reads [149]
@@
[…]
@@ 2.3.1 Global parameters
Droplet size is directly influenced by the amount of droplet material in the system, $\bar{\phi}V_{\mathrm{sys}}$; see equation (2.42). Cells can thus influence whether droplets form and how large they get by producing and degrading the relevant droplet material. Droplet size also depends on the coexisting volume fractions $\phi^{(0)}_{\mathrm{out}}$ and $\phi^{(0)}_{\mathrm{in}}$, which in turn depend on the interaction $\chi$ and the molecular size $\Omega$; see equation (2.21). These aspects can be influenced by modifying the involved molecules, either by post-translational modifications [36, 54] or by multimerisation [172], which can induce phase separation in vitro [173] and in vivo [14, 174]. The interactions between molecules are also sensitive to environmental influences [72], including temperature [175, 176, 177], pH [178], osmotic conditions [179], crowding conditions [69], CO2 concentration [180], and electric fields [181]. Cells can exploit the phase transition associated with changes in these environmental parameters to deploy appropriate responses e.g., to control flowering in plants [182]. There is also evidence that evolution adapts genetic sequences to different temperature niches to control this behaviour [183].
@@ 2.3.2 Spatial gradients
The sensitivity of phase separation to environmental parameters provides exquisite control over droplet positions when external gradients are involved. For instance, compositional gradients can bias droplets toward one side of a system [116, 184, 185, 186, 187], for both passive [150, 188] and active droplets [189], in a process akin to diffusiophoresis [190]. In sufficiently large cells, gravity can also become important, so that droplets sediment due to density mismatch [82]. Finally, artificially induced temperature gradients can position droplets [175], although such gradients are likely negligible in most cells [103].
@@ 3 Beyond simple liquid droplets
Cells contain an endless variety of biomolecules. For example, there are an estimated $10^{5}$ different proteins [199] present in the cell. It is thus not surprising that condensates typically consist of many different types of molecule [200, 201, 202, 55], many of which have complex individual properties themselves. Molecules can also form partially ordered structures within droplets, such as filaments [203] and porous networks [204, 205]. The complex compositional features resulting from this mixture of biomolecules affects both the bulk and interfacial properties of droplets, and also enables the coexistence of multiple different types of droplet within a cell.
@@ 3.1 Complex molecular interactions
Biomolecular condensates contain a huge variety of components, including water, ions, small peptides, proteins and flexible polymers such as RNA and DNA [63, 23]. Since it is challenging to directly study interactions between these molecules in their native environment, much experimental work has thus far relied on in vitro reconstitution [206, 207, 208, 38]. Here, it is crucial to exploit label-free techniques [209, 210] since fluorescent labels typically strongly affect droplet condensation [210, 211]. One key insight from these approaches is that often a few scaffold components control the formation of condensates, while client molecules partition into them [212, 213, 214, 215, 216]. However, this dichotomy is likely an oversimplification and it will be interesting to more rigorously quantify the contributions of individual components, using theoretical analysis [217, 218, 219] or the recently developed dominance metric [220], which can uncover basic interaction motifs between components [221].
It is clear that biomolecular phase separation involves many different kinds of interactions between molecules [222]. To distill the essential features of these interactions, it is often helpful to coarse-grain particular microscopic models [93, 223]. For instance, many biomolecules can be conceptualised as associative polymers [224, 225, 226, 227, 85, 222] and described by sticker-and-spacer models [224, 228, 86, 23, 229], where flexible spacers link stickers with specific interactions. However, real biomolecules are complex and thus require flexible modelling approaches, which are reviewed in refs. [23, 90, 63, 230]. Such models allow elucidating how particular molecular features, like the amino acid sequence of proteins, affect condensate properties [231, 232, 233, 234], and can explain why dilute phases often exhibit surprisingly large clusters [235, 236].
@@ 3.2 Complex material properties
The complex nature of the constituent biomolecules also results in condensates exhibiting complex rheological properties that cannot be captured by simple (Newtonian) droplet models [136], due to long, flexible polymers deforming, entangling, and interacting with each other. Cells exploit these complex material properties to carry out function [243]: for example, centrosomes are able to grow rapidly in a fluid-like manner, but also exhibit solid-like properties to support large stresses during cell division [244, 245].
@@ 3.2.1 Internal stresses
For a single-component Newtonian fluid, the total stress can be decomposed into contributions arising from hydrostatic pressure and viscous dissipation [121]. The hydrostatic pressure $p$ enforces incompressibility of the fluid, and is determined by momentum conservation (Navier–Stokes; equation 2.9) and the incompressibility condition $\nabla\cdot\boldsymbol{v}=0$. Multi-component fluids additionally induce an ‘equilibrium stress’, $\boldsymbol{\sigma}_{\textrm{eq}}$, that captures osmotic and interfacial effects. Further, many biological fluids exhibit a complex rheology that results in an elastic contribution to the stress, denoted by $\boldsymbol{\sigma}_{\textrm{el}}$. The overall stress tensor $\boldsymbol{\sigma}$ then comprises the sum of each of these different contributions,
@@ 3.2.2 Viscoelasticity
Viscoelastic materials can exhibit either liquid-like or solid-like responses to deformation depending on the specific loading conditions. On long timescales (slow deformations), they flow like a liquid, but on shorter timescales (fast deformations), they resist deformation, like an elastic solid. Recent micro-rheological experiments have shown viscoelastic behaviour in a variety of different condensates [136, 250, 251, 252], in which stress relaxation occurs over biologically relevant timescales.
@@
The viscoelastic properties of condensates may change over time, in a process known as ageing [251, 260, 261, 262, 250]. During ageing, condensates progressively become less able to deform, with an increase in relaxation time leading to a more solid-like rheology. The molecular mechanisms that drive condensate ageing are currently an active topic of research [13, 263, 264], although many theoretical works model ageing condensates as soft glassy systems. Aged condensates take longer to recover their shape after deformation and show suppressed ability to coalesce [265]. It is also possible for aged condensates to form complex droplet architectures, in which different parts of the droplet display distinct mechanical properties [266]. Note that the storage and loss moduli of aged condensates at different time points often collapse onto a single curve when scaled by their (time-dependent) characteristic frequency $\omega_{\mathrm{c}}$ [251, 267, 263].
@@ 3.2.3 Gelation
Liquid droplets are characterised by their dynamic, amorphous structure and ability to dissolve in a reversible manner. It is possible, however, for condensates to undergo a phase transition to a gel-like state [14, 268, 242, 269]. This transition results in kinetic arrest, in which condensates maintain their amorphous structure but lose their dynamic properties [270, 271, 272]. Surprisingly, gelation can lead to enhanced ripening relative to liquid droplets [273]. A further transition can result in the formation of irreversible fibrous aggregates, which are linked to the development of various neurodegenerative conditions [28, 13]. Recent work has shown that gelation typically initiates at the interface between the droplet and its surroundings, propagating inwards, such that droplets transiently form a core–shell geometry with a liquid phase surrounded by a gel phase [274]. Similarly, fast changes in composition can result in the formation of kinetically arrested double emulsion structures within droplets[275]. Gelation and aggregation are enhanced in droplets compared to dilute proteins as a result of the increased protein concentration [276, 277], however the formation of metastable non-fibrillar condensates can also suppress aggregation [278]. Gelation can also be driven by shear flows which help to align protein molecules into a kinematically-arrested orientation [279].